Given               B 2 31 2 21 1 32 find the eigenvalues of B and an eigenvector for B corresponding to  1.


1
Expert's answer
2021-09-22T23:57:19-0400
"B=\begin{pmatrix}
 2 & 3 & 1 \\
 2 & 2 & 1 \\
1 & 3 &2
\end{pmatrix}"

"B-\lambda I=\begin{pmatrix}
 2-\lambda & 3 & 1 \\
 2 & 2-\lambda & 1 \\
1 & 3 & 2-\lambda
\end{pmatrix}"

"\begin{vmatrix}
 2-\lambda & 3 & 1 \\
 2 & 2-\lambda & 1 \\
1 & 3 & 2-\lambda
\end{vmatrix}"

"=(2-\lambda)\begin{vmatrix}
 2-\lambda & 1 \\
 3 & 2-\lambda
\end{vmatrix}-3\begin{vmatrix}
 2 & 1 \\
 1 & 2-\lambda
\end{vmatrix}+1\begin{vmatrix}
 2 & 2-\lambda \\
 1 & 3
\end{vmatrix}"

"=(2-\lambda)^3-3(2-\lambda)-6(2-\lambda)+3+6-(2-\lambda)"


"=8-12\lambda+6\lambda^2-\lambda^3-20+10\lambda+9"

"=-\lambda^3+6\lambda^2-2\lambda-3=0"

"-\lambda^2(\lambda-1)+5\lambda(\lambda-1)+3(\lambda-1)=0"

"\lambda_1=1"

"-\lambda^2+5\lambda+3=0"

"D=(5)^2-4(-1)(3)=37"

"\lambda_2=\dfrac{-5+\sqrt{37}}{-2}=-\dfrac{-5+\sqrt{37}}{2}"

"\lambda_3=\dfrac{-5-\sqrt{37}}{-2}=\dfrac{5+\sqrt{37}}{2}"

Eigenvalues: "1, -\dfrac{-5+\sqrt{37}}{2}, \dfrac{5+\sqrt{37}}{2}."


"\lambda=1"



"\begin{pmatrix}
 2-\lambda & 3 & 1 \\
 2 & 2-\lambda & 1 \\
1 & 3 & 2-\lambda
\end{pmatrix}=\begin{pmatrix}
 1 & 3 & 1 \\
 2 & 1 & 1 \\
1 & 3 & 1
\end{pmatrix}"

"R_2=R_2-2R_1"


"\begin{pmatrix}
 1 & 3 & 1 \\
 0 & -5 & -1 \\
1 & 3 & 1
\end{pmatrix}"

"R_3=R_3-R_1"


"\begin{pmatrix}
 1 & 3 & 1 \\
 0 & -5 & -1 \\
0 & 0 & 0
\end{pmatrix}"

"R_2=R_2\/(-5)"


"\begin{pmatrix}
 1 & 3 & 1 \\
 0 & 1 &1/5 \\
0 & 0 & 0
\end{pmatrix}"

"R_1=R_1-3R_2"


"\begin{pmatrix}
 1 & 0 & 2/5 \\
 0 & 1 &1/5 \\
0 & 0 & 0
\end{pmatrix}"

If we take "v_3=t," then "v_1=-\dfrac{2}{5}t, v_2=-\dfrac{1}{5}t,"

Thus


"\vec v=\begin{pmatrix}
 (-2/5)t \\
 (-1/5)t \\
 t
\end{pmatrix}=\begin{pmatrix}
 -2/5)\\
 -1/5 \\
 1
\end{pmatrix}t"

Eigenvalue: "1," eigenvector "\begin{pmatrix}
 -2/5)\\
 -1/5 \\
 1
\end{pmatrix}."